OpenStudy (anonymous):

factor 27^6s+64y^3t 24x^2a-6

OpenStudy (anonymous):

I'm not sure what your question is. Can you use the equation editor to clear it up?

OpenStudy (anonymous):

\[27x ^{6s}+64y ^{3t}\]

OpenStudy (anonymous):

There is no variable that can be factored out so look at the coefficients and see if they have a common factor.

OpenStudy (anonymous):

\[24x ^{2a}-6\]

OpenStudy (anonymous):

In the expression \(27x ^{6s}+64y ^{3t}\) the coefficients are perfect cubes. You can figure that out by prime factorization.

OpenStudy (anonymous):

so i dont know how to do that

OpenStudy (anonymous):

I'm not sure what the question is asking for. Is this on a computer program?

OpenStudy (anonymous):

its asking to factor

OpenStudy (anonymous):

I'm sorry, maybe someone else can help you.

OpenStudy (anonymous):

The first expression looks like it could be the sum of two cubes. [3^(2s)]^3 + [4y^t]^3 Then it could be broken down further, according to a^3 + b^3=(a+b)(a^2 - 2ab + b^2)

OpenStudy (anonymous):

ok but what would i do with the s and t

OpenStudy (anonymous):

@Julian101 wouldn't that mean the s exponent needs to be raised to the third power?

OpenStudy (anonymous):

Nevermind, it doesn't.

OpenStudy (anonymous):

When an exponent is raised to a power, you multiply the exponents.

OpenStudy (anonymous):

but dont you do that when x and y are the same

OpenStudy (anonymous):

its asking to factor

OpenStudy (anonymous):

I'm thinking that "a" would be 3^(2s) and "b" would be 4y^t. So (3^(2s) + 4y^t) (9^(4s) - 2(3^(2s))(4y^t) + 16y^(2t))

OpenStudy (anonymous):

could u write that withe the equation thingie

OpenStudy (anonymous):

I can't seem to get the equation writer to work on my end. It hasn't worked for me in several days, and what the others have written isn't showing up in the equation symbols either, so it's very hard to read.

OpenStudy (anonymous):

\((3x^{2s} + 4y^t) (9x^{4s} - 2(3^{2s})(4y^t) + 16y^{2t})\) I think this is Julian's expression.

OpenStudy (anonymous):

then would that be like the final answer

OpenStudy (anonymous):

I added an x after the 3 and 9, I should have added an x after the 3 in the middle term as well.

OpenStudy (anonymous):

\((3x^{2s} + 4y^t) (9x^{4s} - 2(3x^{2s})(4y^t) + 16y^{2t})\)

OpenStudy (anonymous):

As to that being the final answer, I don't know. Maybe those middle terms need to be multiplied.

OpenStudy (anonymous):

I think I made a mistake, the 2 should be removed from the middle of the second expression in brackets.

OpenStudy (anonymous):

no there is something wrong. \[a ^{3}+b ^{3}=\left( a+b \right)\left( a ^{2}-ab+b ^{2} \right)\]

OpenStudy (anonymous):

a^3 +b^3 = (a+b) (a^2 -ab +b^2) i.e., there is NOT a 2 in front of the ab

OpenStudy (anonymous):

Sorry about that. Other than that, I think that should be final, as I don't think it can be reduced further.

OpenStudy (anonymous):

For the second question, factor out 6, so now you have 6(4x^2a -1) then break down the expression in brackets = 6(2x^a - 1)(2x^a + 1)